OpenStudy (anonymous):

Use De Moivre's theorem to write the complex number in trigonometric form: (cos(3pi/12) + i sin (3pi/12))^5 A) cos(15 pi/12) + isin (15pi/12) B) 5(cos(3 pi/12) + isin (3 pi/12) C) Cos (3pi/60) + isin (3pi/60) D) Cos (3pi/12)^5+i sin (3pi/12)^5

OpenStudy (anonymous):

What is De Moivre's theorem?

zepdrix (zepdrix):

De Moivre's Theorem let's us deal with the exponent on the outside: \[\Large\rm \left[\cos\left(\color{orangered}{\theta}\right)+\mathcal i \sin\left(\color{orangered}{\theta}\right)\right]^c=\left[\cos\left(\color{orangered}{c\theta}\right)+\mathcal i \sin\left(\color{orangered}{c\theta}\right)\right]\]So just bring it inside as a coefficient multiplying your angle.

OpenStudy (anonymous):

I feel like in that case it'd be either C or D, right?!

zepdrix (zepdrix):

\[\large\rm \left[\cos\left(\color{orangered}{\frac{3\pi}{12}}\right)+\mathcal i \sin\left(\color{orangered}{\frac{3\pi}{12}}\right)\right]^5=\left[\cos\left(\color{orangered}{5\cdot\frac{3\pi}{12}}\right)+\mathcal i \sin\left(\color{orangered}{5\cdot\frac{3\pi}{12}}\right)\right]\]

zepdrix (zepdrix):

Remember how to multiply a whole number by a fraction? Just bring it up into the numerator.

OpenStudy (anonymous):

got it! so in that case it'd be A, right?!

zepdrix (zepdrix):

Mmmm yes, good job c:

OpenStudy (anonymous):

thank you for your help!!!